Presentation is loading. Please wait.

Presentation is loading. Please wait.

June 22, 2007Hall A Collaboration Meeting Hall A Beamline Activation John J. LeRose.

Similar presentations


Presentation on theme: "June 22, 2007Hall A Collaboration Meeting Hall A Beamline Activation John J. LeRose."— Presentation transcript:

1 June 22, 2007Hall A Collaboration Meeting Hall A Beamline Activation John J. LeRose

2 June 22, 2007Hall A Collaboration Meeting What happened? March 26, 2007: High Radiation levels are found in Hall A after what was thought to be a relatively quiet weekend of running.

3 June 22, 2007Hall A Collaboration Meeting

4 June 22, 2007Hall A Collaboration Meeting Thursday March 22 Experiment is running smoothly. Two partial surveys indicate no excessive activation in the Hall.

5 June 22, 2007Hall A Collaboration Meeting Friday March 23 –MCC takes the beam away during day shift to work on Hall Bs beam quality. Attempts to fix things solely in the fifth pass fail and adjustments are made as far back as 2nd pass as well. –After cleaning up the Hall B beam MCC has trouble finding a match for Hall A. –16:18 40 μA CW beam restored to the Hall. Radiation monitor in the tunnel just upstream of the Hall, RM-29, immediately shows a 3 order of magnitude increase in γ-radiation compared to Thursday. Neutron radiation at RM-29 goes up markedly as well. This condition remains for the rest of the weekend. Activity in the target ion chambers is up but not significantly. Compton and ep ion chambers are quiet. –Experiment runs smoothly through the night with minor concerns re possibly partially hitting the target frame with the rastered beam.

6 June 22, 2007Hall A Collaboration Meeting Saturday March 24 –Owl shift runs smoothly –10:06 harps scans indicate beam is OK (small spread, y~150 μm) –11:30 MCC takes beam to work on energy spread (30 minutes) –Afternoon has repeated attempts to work on the energy spread. 1C12 OTR indicates a tail on the beam. Harp scans indicate small energy spread but SLI disagrees. Ion chambers are still relatively quiet. RM-29 continues to scream. –Swing shift: We have good shift. Not any problem. Log is full of discussion of raster and target frame, but no one is really concerned.

7 June 22, 2007Hall A Collaboration Meeting Sunday March 25 –More smooth sailing –MCC takes beam for 5 minutes at 13:30 to investigate an orbit problem. –Compton and ep ion chambers show a small peak in activity mid-day, but well below any alarm thresholds. –Swing shift is concerned about rastered beam hitting the target frame but continues to take data. –18:25 MCC makes an adjustment of the vertical beam spotsize (238 μm 113 μm)

8 June 22, 2007Hall A Collaboration Meeting Monday March 26 –01:31 target camera fails –04:30 Hall access to reset right NMR. Remote reset doesnt work anymore. Partial survey indicates some elevated radiation levels. Only whole body doses were recorded on the path to the right NMR. In keeping with ALARA no one lingered to make a full survey. –09:15 Full survey of the Hall indicates serious activation all along the beamline.

9 June 22, 2007Hall A Collaboration Meeting General Comments/Observations Hall A Shiftworkers were unaware that there was any potentially activation causing problem with the beam. –Concern re rastered beam striking target frame existed the entire run and doesnt explain activation upstream The spot became increasingly wide over the weekend ( μm). Singles rates never varied by more than 10%. The target OTR was not working –Requested by experimenters in advance –Might have been useful

10 June 22, 2007Hall A Collaboration Meeting General Comments/Observations MCC On the Accelerator side there was a similar lack of realization that there was any serious problem. Work done on Friday aimed at meeting Hall Bs specs involved making adjustments all the way back to 2nd pass. After that there was difficulty finding a match for Hall A. –That was the case for the whole running period. Knew the beam was not as good as it could be but problem indicators like the EP, Compton, and Target Ion Chambers showed some extra activity but nothing alarming. Detailed harp scans performed on Saturday at ~5 μA in fact show pretty good beam with very small energy spread (σE ~ 2x10-5). –However, there was difficulty getting the SLI at 1C12 to show the same small energy spread. – Arne Freyberger has indicated that he thinks this is because of a beam loading problem that sets in above 10 μA Beam loss monitors also indicated no significant problems Target Ion Chambers Mar (R/hr). Trip level is at 3000 EP Ion Chambers Mar (R/hr). Some increased low level activity mid-day Sunday. Spike mid-day Saturday is from harp scans Compton Ion Chambers Mar (R/hr), similar to EP.

11 June 22, 2007Hall A Collaboration Meeting Radiation monitor output for RM-29 located in the Hall A tunnel. (γs on the left, ns on the right). Both γs and ns, especially γs, show a dramatic increase during the weekend in question.

12 June 22, 2007Hall A Collaboration Meeting Conclusions Indicators on both the accelerator and experimental sides, with the notable exception of RM-29, indicated that the beam was acceptable. All of our alarms and interlocks are geared toward sudden dramatic failures in beam transport. Radiation monitor readings suggest that the beam, probably a low energy tail on the beam, scraped against something in the Hall A tunnel just past the end of the arc (near or at the green wall) producing a cascade that tumbled down the beamline activating things along the way. –First suggested by Pavel Degtiarenko In the survey of the morning of March 26 one sees that radiation levels on the right side of the beamline are uniformly slightly higher than on the left, supporting the low energy tail hypothesis.

13 June 22, 2007Hall A Collaboration Meeting Recommendations Shift workers should observe the radiation monitor levels, RM-29 in particular, and alert MCC, the run coordinator, and experimental spokespersons when high radiation levels are observed (> 5R/hr γs at RM-29) for extended periods of time (>10 minutes). This is a very good indicator that beam is not coming cleanly into the Hall Machine Ops should setup something similar. If they have a strong philosophical objection to using RADCON instruments in the machine protection scheme, then at least some sort of integrated ion chamber signal separate from the peak level interlocks could be used. Although, it appears the ion chambers are less sensitive to this kind of lower level radiation (Ion chambers report in increments of 10 R/hr )

14 June 22, 2007Hall A Collaboration Meeting Compton Ion Chambers Mar (R/hr), similar to EP.

15 June 22, 2007Hall A Collaboration Meeting EP Ion Chambers Mar (R/hr). Some increased low level activity mid-day Sunday. Spike mid-day Saturday is from harp scans

16 June 22, 2007Hall A Collaboration Meeting Target Ion Chambers Mar (R/hr). Trip level is at 3000


Download ppt "June 22, 2007Hall A Collaboration Meeting Hall A Beamline Activation John J. LeRose."

Similar presentations


Ads by Google